OpenStudy (anonymous):

use the method of subsitution 7a=2-b 7a=6-b

OpenStudy (anonymous):

I would subistute a=x and b=y 7x=2-y 7x=6-y y=7x-2(6-7x) 7x-12-21x Right?

OpenStudy (anonymous):

excpet that it would be 7x+12-21x

hartnn (hartnn):

i would have substituted 7a = 2-b from 1st equation into 2nd equation to get 2-b=6-b

OpenStudy (anonymous):

@hba I need help again I know this is similar to the one that I so easily got confused with but this has a and b can you help me???

OpenStudy (anonymous):

@hartnn I got confused at what you said how would I put it with the second equation, 7a=6-b(7a=2-b)?

hartnn (hartnn):